Subject: APO 3.5m users committee minutes, December 2010
From: Suzanne Hawley
Submitted: Fri, 10 Dec 2010 19:41:22 -0800 (PST)
Message number: 1212
(previous: 1211,
next: 1213
up: Index)
APO 3.5-m Users Committee Phonecon, 12/06/10
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
Russet McMillan, Remy Indebetouw, Scott Anderson, John Bally, Al Harper,
Sean Moran, Jon Holtzman, Bruce Gillespie
Absent: None
**********************************
User feedback, comments from institutional representatives:
o Chicago (Al Harper) - Al had nothing to report.
o New Mexico State (Jon Holtzman) - Jon had nothing to report.
o Washington (Scott Anderson) - Scott had nothing to report.
o Colorado (John Bally) - John has heard no comments except from one
user regarding the recent problem with the secondary motion controls.
o Princeton (Michael Strauss) - Michael said the PU users are happy,
no complaints. Their one NIC-FPS user has no plans for using the
instrument in the upcoming quarter.
o Johns Hopkins (Sean Moran) - Sean had nothing to report.
o Virginia (Remy Indebetouw) - Remy had nothing to report.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark made note of the recent spate of good observing weather at APO.
He highlighted the reexamination of humidity closure limits and
humidity-sensing equipment upgrades, which have led to a recent
change in closure limits to allow a slightly smaller dew point
depression (but only when the humidity is stable). In addition, new
humidity-sensing equipment is being obtained that should enable us to
stay open even closer to the dew point. After some discussion, there
seemed to be a consensus that the amount of time we are closed due to
humidity under clear skies is a few percent of the year, and is
especially concentrated in the summer monsoon period. The new
procedures and sensing equipment should help reduce this somewhat.
The Agile instrument rotator problem was tracked down and fixed.
DIS is performing well, but we've recently had a recurrence of the
sticky slit-view camera shutter when cold -- we put in the spare shutter
and are buying a replacement. There was some condensation on the
SPIcam detector, which disappeared after pumping on the dewar. Recent
incidents of slit moves "not reporting" in TripleSpec were determined
to be due to the slit not moving when commanded. Apparently this occurred
for most of November, and was recently fixed. All affected Triplespec
users were notified. There has been intermittent low-level banding in Echelle
data. According to users who have reduced recent echelle data, the
effects are removed during data reduction and are not affecting the final
spectra. However, we are discussing the situation with Dale Sandford
at Yerkes, who is planning a trip to APO next quarter to try and
diagnose the problem. Finally, when NIC-FPS is rotated to angles
greater than about 90 degrees, the data often becomes corrupted.
When this happens the instrument can be reset to allow data collection
to continue, but the problem often returns. The external electronics
are being investigated as a possible cause, but we suspect that the
problem is inside the dewar.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 10/6/10
through 12/01/10
0) Overview
Lots of clear and dry weather. We have adjusted the humidity opening
and closing limits slightly to provide more on-sky time with further
improvements possible once new metrology is obtained and operational.
Major issue this period has been NIC-FPS readout when rotated upside
down. See below for more details.
1) Telescope
The telescope was re-collimated by adjusting M1 and M2 positions allowing
for a more nominal M2 position at the center of range. Agile rotator
amplifier is occasionally going in to a low-torque mode. We found a
ground loop problem, and while more testing is needed the problem has
not reoccurred since this was fixed. We fought a lot of Miller moth
debris on the drive discs towards the end of the moth season. Daily
cleanings helped, but several times cleanings during the night were
required. TCC version 2.14.0 was also implemented.
2) Instruments
DIS is having cold temperature issues with the slitviewer shutter.
A replacement shutter was reinstalled but is still having some
intermittent failures when cold.
SPICam was discovered to have some condensation on the detector. After
vacuum pumping, it appears to all be gone. Issue with exposure times
was resolved (see later agenda item).
The Agile rotator still occasionally faults out due to insufficient
torque. Ed discovered a ground loop and is currently running the
rotator during the day to see if the problem repeats. A change to
the umbilical retractor was also made to reduce the load on the rotator.
TripleSpec appears to not be reporting slit moves correctly. They are
reported correctly with a status but not during a move. We have asked
the instrument team to look at this.
Echelle started showing some banding in bias images. They show up
intermittently and are now under investigation.
NIC-FPS images often do not readout when the instrument is rotated past 90
degrees. Much testing has been done on all "outside the dewar" components
and nothing has helped. We have pretty much isolated the problem to inside
the dewar and are waiting for a response from CU on how to proceed. The
instrument seems to be OK with limited rotation and recovery is quick
although not always 100% effective. NIC-FPS is only available in a
shared risk state at this time.
**********************************
Instrument issues - Hawley:
Suzanne, Jon, and Russet talked about some of the recent instrument
issues and topics. It was recently confirmed that the SPIcam exposure
times are not being correctly recorded -- the commanded and actual
exposure times differ by 3/4 second; this has been the case since the
summer of 2007. Craig and Fritz found the source of the problem, which
occurred during the conversion to TUI, and it has been fixed as of the
week of 11/22. Commanded and actual exposure times are now the same,
and users who took SPIcam data in the '07 through '10 interval have been
notified. Russet has double-checked that the correct exposure times
are now getting written in the data headers.
For NIC-FPS, we have the ongoing problem of high read noise, and the new
fan-out board has yet to be installed. More recently, we have the unsolved
problem of readout errors at upside-down rotation angles. Taken together,
the NIC-FPS users need to be cognizant of these issues, and if they use
the instrument it is on a shared-risk basis. The ARC Board of Governors
were apprised of the current situation last month, and they endorsed the
plan of taking the instrument out of service and sending it back to
CU for repairs.
The recent cluster of instrument problems shows the effects of equipment
aging, and the need for next-generation instrumentation. We need ARC
astronomers to come forward and lead initiatives for new instruments
for the 3.5-m telescope.
**********************************
Board meeting, futures discussion - Hawley:
The ARC Board of Governors meeting went well, and the Board members
were particularly impressed with the small amount of unscheduled down
time in the past year (~21 hours). The flexibility that the facility
afforded the users was also appreciated. The 2011 budget was unanimously
approved. On the ARC Futures topic, four pre-proposals were received
that involve the 3.5-meter, and nearly double that number were received
for future 2.5-meter programs. Suzanne will convene a review team to
look at the 3.5-meter pre-proposals to help decide how to proceed with them.
**********************************
RestAR meeting, upcoming proposal deadline - Holtzman, Hawley:
Jon attended the recent ReSTAR meeting in Tucson. There is NOAO interest
in obtaining access to mid-sized telescopes for use by the national
community. In the previous round of ReSTAR funding, arrangements were
made for access to the Palomar 5-meter telescope, and Cornell was engaged
to build a copy of TripleSpec for use by the national observatories. In
this next phase of ReSTAR, NOAO is looking for proposals (due in a couple
of months) from facilities that wish to team with NOAO to write a joint
proposal to the NSF. Suzanne said that we've been musing for some time
about selling observing time to NOAO if we needed funding, but it is
not obvious that we need to pursue this path at this time. We can also
imagine other long-term arrangements with NOAO that might be outside
of the ReSTAR process -- this will be discussed further.
**********************************
Guider images attached to spectra headers - McMillan, Bally:
John Bally offered the idea of recording guide images in the headers
(or FITS extensions) for DIS spectra. Russet pointed out that TripleSpec
already writes a slit-view guider image with the spectrum. She felt
that we could also do this for DIS and possibly Echelle. John added
that he would be interested in the slit-viewer images, plus various
keywords (such as guider status, etc.). Suzanne said that the APO
staff will look into what is required to implement this for DIS.
**********************************
Other business - group:
Scott congratulated Suzanne for accepting the ARC Board's offer for a
third term as the 3.5-meter Director.
Jon said that he's been working on the echelle upgrade project, and
he recently pinged the instrument development group at JHU for their
assessments on some of the proposed upgrades.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Users Committee members should poll their users to find
out how many are using TUI with Mac OS 10.4 and report at the next Users
Committee meeting.
STATUS: No affected users were identified; this action is CLOSED.
[new actions from this meeting]:
None.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 31
January at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1212 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO